### v3

# Fault Tolerant Zero Forcing

Abstract
Zero forcing is an iterative graph coloring process studied for its wide array of applications. In this process, the vertices of the graph are initially designated as blue or white, and a zero forcing set is a set of initially blue vertices that results in all vertices becoming blue after repeated application of a color change rule. The zero forcing number of a graph is the minimum cardinality of a zero forcing set. The zero forcing number has motivated the introduction of a host of variants motivated by linear-algebraic or graph-theoretic contexts. We define a variant we term the $k$-fault tolerant zero forcing number, which is the minimum cardinality of a set $B$ such that every subset of $B$ of cardinality |B|k|B|-k is a zero forcing set. We study the values of this parameter on various graph families, the behavior under several graph operations, and characterize the 1-fault tolerant zero forcing number of trees.
Keywords: zero forcing, fault tolerant, graph operations, path cover
MSC Classification: 05C50, 05C69
1 Introduction
Zero forcing on a graph was introduced in [2] to bound the maximum nullity obtained in space of matrices defined by the graph. We delay formal definitions until Section 2, but informally the goal of zero forcing is to determine the smallest size of a set of vertices that when these vertices are the only vertices initially colored blue, the remaining vertices all become blue after a finite number of applications of the zero forcing color change rule.
Since its introduction, there has been significant interest in zero forcing and a host of related parameters that are defined using a similar paradigm. Though first studied for its applications to the problem of determining the minimum rank of a symmetric matrix associated with a graph [2, 5, 8, 15], zero forcing and its variants have also been studied for their applications to quantum control [11], the spread of disease and information [13], and monitoring electric grids[19]. The latter application is related to the placement of phase-monitoring units (PMUs) in an electrical power network, and thus it has motivated investigation into variations of the parameter when PMUs are allowed to fail [9, 25].
The property of robustness when searching a graph is a desirable property, and in this article we introduce a fault tolerant variant of zero forcing and study its properties. While the leaky-forcing variant of zero forcing studied in [1, 3, 4] seeks to provide robustness of the zero forcing process when certain vertices are not allowed to contribute to the zero forcing process, our parameter is determined by the behavior of all subsets of a certain size within a prescribed set of vertices. Thus our parameter is novel and distinct.
In [23], zero forcing is related to the solving of linear systems, and leaky forcing was defined as a parameter to study the resilience of solving linear or physical systems where a leak at a node prevents the information available at that from being able to be used reliably in deducing state of other vertices in the system. In contrast, the fault tolerant zero forcing parameter studied herein corresponds to a loss of information at a node (or nodes) but not the reliability of information about the node through the rest of the system.
Much is known about the zero forcing number, including the values of the parameter for many families of graphs, as well as its behavior under various graphs operations, see [21] for a detailed account. Additionally, other parameters of a graph derived directly from the zero forcing number and the zero forcing color change rule have been investigated. Examples include the propagation time [20] of a graph and the throttling number [12] of a graph. We thus have many tools available for studying the fault tolerant zero forcing parameter.
The remainder of the article is organized as follows. We give the foundational definitions in Section 2, as well as make a number of observations useful to the study of fault tolerant zero forcing. Section 3 is devoted to determining the fault tolerant zero forcing number for several common graph families. A general lower bound on the fault tolerant zero forcing number is given in Section 4, where we also characterize the fault tolerant zero forcing number for trees. Finally, the effect of various graph operations on the fault tolerant zero forcing number is explored in Section 5.
2 Preliminaries
Let $G$ be a simple graph (i.e. undirected with no loops or multiple edges) and denote its set of vertices by $V(G)$ and its edges by $E(G)$. We denote by $uv$ the edge of $G$ whose endpoints are the vertices $u,v\in V(G)$, while $\{u,v\}$ denotes the subset of $V(G)$ with these vertices. For $V\subseteq V(G)$, we denote the subgraph of $G$ induced by $V$ by $G[V]$. We follow the notation in [21], and in particular for a graph with $n$ vertices, we denote path graphs by $P_{n}$, cycle graphs by $C_{n}$, and complete graphs by $K_{n}$. The complete bipartite graph with partite sets of $m$ and $n$ is denoted $K_{m,n}$. We denote the degree of a vertex $v\in V(G)$ by $\deg_{G}(v)$ or $\deg(v)$ if the context is clear. A vertex $v$ is called universal if $\deg(v)=n-1$, and $v$ is called a leaf if $\deg(v)=1$. Wheels graphs $W_{n}$ are constructed by adding a universal vertex to $C_{n-1}$. The symbols $N(v)$ and $N[v]$ denote the open and closed neighborhoods of $v$, respectively. Two vertices $u$ and $v$ of $G$ are twins if $N_{G}(u)=N_{G}(v)$.
We choose an initial set of vertices $B\subseteq V(G)$ to designate as blue, and the complement of $B$ in $V(G)$, denoted $W$, are designated white. The zero forcing color change rule allows a blue vertex $b\in B$ to change the color of a white vertex $w\in W$ to blue if $w$ is the unique white neighbor of $b$. In this case, we say that $b$ forces $w$ and this is denoted by $b\rightarrow w$. The zero forcing color change rule is applied iteratively until no more forces are possible. We say that the initial set of blue vertices $B$ is a zero forcing set for $G$ if eventually all of $V(G)$ is colored blue under repeated application of the color change rule, and $B$ is a failed zero forcing set otherwise. An example is shown in Figure 2.1.
11223344 55 (a)11 22 33 44 55 $(b)$
Figure 2.1: A minimum zero forcing set for $G$ in $(a)$, and a failed zero forcing set for $G$ in $(b)$.
Clearly, one could choose $B=V(G)$ to obtain a zero forcing set, so the interesting cases are when $B$ contains as few vertices as possible. This leads to the definition of $\operatorname{Z}(G)$, the zero forcing number of a graph $G$.
Definition 2.1.
Let $G$ be a graph. The zero forcing number of $G$, denoted $\operatorname{Z}(G)$, is the minimum cardinality of a zero forcing set for $G$.
A zero forcing set can be thought of as a set of vertices such that if that status of the vertices in the set are known, then the status of every vertex in the graph is determined. For instance, in a space of matrices described by the graph, if the values of entries in a null vector corresponding to a zero forcing set are chosen to be zero, the null vector must be the zero vector. See [21] for full details.
By introducing a robustness requirement, as in our next definition, we ensure that a change in status of one or more of the vertices in a zero forcing set does not change the deterministic nature of the set.
Definition 2.2.
Let $G$ be a graph and $B$ a subset of the vertices $V(G)$. Suppose $B$ has $m$ elements, i.e. $|B|=m$. Then $B$ is a fault tolerant zero forcing set if every subset $F\subset B$ such that |F|=m1|F|=m-1 is also a zero forcing set. Similarly, $B$ is $k$-fault tolerant if every subset $F\subset B$ such that |F|=mk|F|=m-k is also a zero forcing set.
Definition 2.3.
Let $G$ be a graph. The $k$-fault tolerant zero forcing number of $G$ is the minimum cardinality of a $k$-fault tolerant zero forcing set for $G$. i.e.
$\operatorname{Z}_{t}^{k}(G)=\min\left\{|B|\text{ such that }B\text{ is $k$-fault tolerant}\right\}$
When $k=1$, we will use $\operatorname{Z}_{t}(G)$ to denote $\operatorname{Z}_{t}^{1}(G)$. Let $B$ be a $k$-fault tolerant set and $S\subset B$ such that $|S|=k$. When discussing the forcing process of $B\setminus S$, we refer to $v\in S$ as a fault.
In the remainder of this section, we will make some simple observations about the fault tolerant zero forcing number. Though each is quite straight forward, they are nevertheless useful in later sections.
First, we observe that the fault tolerant zero forcing number operates independently over connected components of graphs.
Observation 2.4.
If $G$ is a graph, its connected components are $G_{1},G_{2},\dots,G_{r}$, and $\operatorname{Z}_{t}^{k}(G_{i})$ exists for $i=1,2,\dots,r$, then
$\operatorname{Z}_{t}^{k}(G)=\displaystyle\sum_{i=1}^{r}\operatorname{Z}_{t}^{k}(G_{i}).$
In light of this observation, there is no loss of generality in considering only connected graphs. It is natural to look for bounds on the fault tolerant zero forcing number $\operatorname{Z}_{t}^{k}(G)$ in terms of the zero forcing number $\operatorname{Z}(G)$, and our next observation establishes such a bound.
Observation 2.5.
Let $G$ be a graph and suppose $\operatorname{Z}_{t}^{k}(G)$ exists. Then $\operatorname{Z}_{t}^{k}(G)\geq\operatorname{Z}(G)+k$. Moreover,
$\operatorname{Z}_{t}^{k}(G)>\operatorname{Z}_{t}^{k-1}(G)>\dots>\operatorname{Z}_{t}(G)>\operatorname{Z}(G).$
We will show in Section 3 that several common graph families, such as paths, stars, and complete graphs, have equality with this lower bound when $k=1$. Note that these families all have extreme zero forcing numbers relative to their order. When the zero forcing number of a graph is lower, such as cycles and wheels, inequality is obtained. See Section 3 for more details.
If a graph $G$ contains two minimum zero forcing sets with no vertices in common, then choosing the union of those sets provides a fault tolerant zero forcing set and establishes that $\operatorname{Z}_{t}(G)\leq 2\operatorname{Z}(G)$, but in practice this is not a very useful bound.
Definition 2.6.
Let $G$ be a graph and $B$ a zero forcing set of $G$. We say $B$ is a minimal zero forcing set if for all $v\in B$, $B\setminus\{v\}$ is not a zero forcing set of $G$.
An important distinction exists between minimum zero forcing sets and minimal zero forcing sets. To illustrate the difference, we consider the path on $8$ vertices, labeled as in Figure 2.2. There are exactly two minimum zero forcing sets, namely $\{v_{1}\}$ and $\{v_{7}\}$, and these sets are also minimal. There are four additional minimal zero forcing sets of $G$, namely $\{v_{2},v_{3}\}$, $\{v_{3},v_{4}\}$, $\{v_{4},v_{5}\}$, and $\{v_{5},v_{6}\}$.
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$v_{7}$
Figure 2.2: The graph $P_{7}$.
Every fault tolerant zero forcing set contains a minimal zero forcing set, but not necessarily a minimum zero forcing set, as the next example demonstrates.
Example 2.7.
115533662244
Figure 2.3: The graph in Example 2.7.
Consider the graph Figure 2.3. The minimum zero forcing sets are $\{1,3\},\ \{2,4\},\{3,5\},$ and $\{4,6\}$ The set of vertices $B=\{1,2,5,6\}$ is a minimum fault tolerant set, but $B$ does not contain any of the minimum zero forcing sets.
The study of zero forcing and its variants has been augmented by the introduction of the notion of forts, which are obstructions to zero forcing sets. Specifically, a fort is a nonempty subset $F\subseteq V(G)$ such that for all $v\notin F,|N_{G}(v)\cap F|\neq 1$. Intuitively, a fort is a set of unfilled vertices such that when its complement if filled, no more forces are possible.
Theorem 2.8 ([10],[17]).
Let $G$ be a graph. A set of vertices $S\subseteq V(G)$ is a zero forcing set if and only if $S$ intersects every fort of $G$ nontrivially.
Theorem 2.8 extends naturally to the vertex fault tolerant case.
Observation 2.9.
A set of vertices $S\subseteq V(G)$ is a fault tolerant zero forcing set if and only if $S\cap F\geq k$ for every fort $F$ of $G$.
Proposition 2.10.
Let $G$ be a graph. Then:
  1. 1.
    The only forts of $G$ of cardinality 1 are isolated vertices.
  2. 2.
    The forts of $G$ of cardinality 2 are precisely the sets of (adjacent or independent) twins.
Proof.
Item 1 follows from the fact if $\{v\}$ is a fort, then any neighbor $u$ of $v$ satisfies $|N_{G}(u)\cap\{v\}|=1$, and therefore $v$ has no neighbors. For Item 2, it is clear that a pair of twins is a fort of cardinality 2. If $\{u,v\}$ is fort, then $N_{G}(u)=N_{G}(v)$ since any vertex $w$ in one neighborhood but not the other would satisfy $|N_{G}(w)\cap\{u,v\}|=1$. ∎
For every graph $G$, the standard zero forcing number $\operatorname{Z}(G)$ always exists. However, when fault tolerance is enforced, the existence of $\operatorname{Z}_{t}^{k}(G)$ is not guaranteed. Proposition 2.10 and 2.9 allow us to characterize the existence of $\operatorname{Z}_{t}^{k}(G)$ for $k\leq 2$.
Theorem 2.11.
Let $G$ be a graph. Then:
  1. 1.
    $\operatorname{Z}_{t}(G)$ exists if and only if $G$ has no isolated vertices.
  2. 2.
    $\operatorname{Z}_{t}^{2}(G)$ exists if and only if $G$ has neither isolated vertices nor twins.
  3. 3.
    If $u,v$ are twins of $G$, then $\{u,v\}$ is a subset of every fault tolerant zero forcing set.
Note that in particular, Theorem 2.11 implies that if $k\geq 2$, then $\operatorname{Z}_{t}^{k}(G)$ does not exist if $G$ is a complete graph or complete bipartite graph.
3 Graph Families
We now establish values of the fault tolerant zero forcing number for graphs which belong to a common family of graphs. The values of the zero forcing number for these graph families were established in [2].
Proposition 3.1.
Let $G=P_{n}$ be a path on $n\geq 2$ vertices. Then $\operatorname{Z}_{t}(G)=2$. If $n\geq 4$, then $\operatorname{Z}_{t}^{2}(G)=4$.
Proof.
Let $G$ be a path on $n\geq 2$ vertices. Then $\operatorname{Z}(G)=1$ and by the 2.5 we have $\operatorname{Z}_{t}(G)\geq\operatorname{Z}(G)+1=2$. By choosing $B$ to consist of both endpoints, we are guaranteed that every subset of size 1 of $B$ is a zero forcing set, and thus $B$ is a fault tolerant zero forcing set. Since this achieves the lower bound, we have $\operatorname{Z}_{t}(G)=2$.
Now suppose $n\geq 4$. The graph $G$ contains no 2-fault tolerant sets of size 3, since if the two vertices of degree 1 are white there are no zero forcing sets of size 1. Clearly if $B$ consists of a pair of adjacent degree 2 vertices and the degree 1 vertices, then $B$ is a 2-fault tolerant set since it contains 3 disjoint zero forcing sets. ∎
Proposition 3.2.
If $G=K_{n}$ for $n\geq 2$, then $\operatorname{Z}_{t}(G)=n$.
Proof.
Let $G=K_{n}$ for some $n\geq 2$. Then $\operatorname{Z}(K_{n})=n-1$ for all $n\geq 2$, and any set of n1n-1 vertices is a minimum zero forcing set. Therefore, the set of all $n$ vertices is a fault tolerant zero forcing set, and thus $\operatorname{Z}_{t}(K_{n})\leq n$. We have $\operatorname{Z}_{t}(K_{n})\geq n$ by 2.5. ∎
Proposition 3.3.
If $G=K_{1,n}$ for $n\geq 3$, then $\operatorname{Z}_{t}(G)=n$.
Proof.
$\operatorname{Z}(G)=n-1$, and all minimum zero forcing sets are composed of all but one leaf. Thus, the set of all leaves is a 1-fault tolerant set, so $\operatorname{Z}_{t}(G)\leq n$. We have $\operatorname{Z}_{t}(G)\geq n$ by 2.5. ∎
Proposition 3.4.
If $G=K_{m,n}$ for $n\geq m\geq 2$, then $\operatorname{Z}_{t}(G)=m+n$.
Proof.
Let $G=K_{m},n$ for some $n\geq m\geq 2$. The set $B=V(G)$ is a fault tolerant zero forcing set of $G$ since only one force is required for each $B\setminus\{v\}$, so $\operatorname{Z}_{t}(G)\leq m+n$.
Suppose there is a smaller fault tolerant zero forcing set $S$ of $G$. Denote the partite sets of $G$ by $A$ and $B$, then coloring the vertices of $S$ blue leaves at least one white vertex $v$ initially, assume without loss of generality $v\in A$. There is always a vertex $w\in S$ such there are two white vertices $v,p\in A$ after setting the vertices of $S\setminus\{w\}$ blue. Since $N_{G}(v)=N_{G}(p)$, neither $v$ now $p$ can be forced and $S\setminus\{w\}$ is a failed zero forcing set. ∎
Proposition 3.5.
If $G=C_{n}$ for $n\geq 4$, then $\operatorname{Z}_{t}(G)=4$.
Proof.
The minimum (and minimal) zero forcing sets of $C_{n}$ consist of pairs of adjacent vertices, thus $\operatorname{Z}(C_{n})=2$ for $n\geq 3$. By 2.5, we have that $\operatorname{Z}_{t}(C_{n})\geq 3$. However, no set of three vertices guarantees that two adjacent vertices remain after one is removed. A path of length 4 in $C_{n}$ preserves this property, and this is only possible if $n\geq 4$. ∎
Proposition 3.6.
If $G=W_{n}$ is a wheel on $n\geq 6$ vertices, then $\operatorname{Z}_{t}(G)=5$.
Proof.
Let $G=W_{n}$ for $n\geq 6$. It was shown in [16] that $\operatorname{Z}(G)=3$. Furthermore, up to isomorphism there are two minimum zero forcing sets, and these are precisely the minimal zero forcing sets. The first minimum zero forcing set is $B_{1}=\{v_{1},v_{2},v_{3}\}$, where $\deg(v_{1})=\deg(v_{2})=3$, $v_{1}v_{2}\in E(G)$, and $\deg(v_{3})=n-1$, while the second is $B_{2}=\{w_{1},w_{2},w_{3}\}$, where $\deg(w_{1})=\deg(w_{2})=\deg(w_{3})=3$ and $w_{1}w_{2},w_{2}w_{3}\in E(G)$. By 2.5, we have $\operatorname{Z}_{t}(G)\geq 4$. We will show there is no $B$ such that $|B|=4$ is a fault tolerant zero forcing set and construct a fault tolerant zero forcing set $B$ such that $|B|=5$.
Suppose $B$ is a fault tolerant zero forcing set of $G$ and $|B|=4$. Then $B$ must contain either $B_{1}$ or $B_{2}$ as a subset. Suppose $B=B_{1}\cup\{v_{4}\}$. If $v_{4}\not\in N(\{{v_{1},v_{2}}\})$, then $B\setminus\{v_{1}\}=\{v_{2},v_{3},v_{4}\}$ and each of these vertices has two white neighbors, and thus no forces are possible. If $v_{4}\in N(\{{v_{1},v_{2}}\})$, then assume without loss of generality that $v_{1}v_{4}\in E(G)$. Again, every vertex of $B\setminus\{v_{1}\}=\{{v_{2},v_{3},v_{4}}\}$ has two white neighbors and no forces are possible. Therefore, there are no fault tolerant zero forcing sets of cardinality four that contain $B_{1}$.
Now suppose $B=B_{2}\cup\{w_{4}\}$. If $w_{4}$ is the universal vertex, we have already shown that $B$ is not fault tolerant. If $w_{i}w_{4}\not\in E(G)$ for $i=1,2,3$, then no forces are possible from $B\setminus\{w_{1}\}$, so without loss of generality assume $w_{3}w_{4}\in E(G)$. In this case $B\setminus\{w_{2}\}$ is not a zero forcing set, so $B$ cannot be fault tolerant. Thus $\operatorname{Z}_{t}(G)\geq 5$.
Now suppose $B=\{{v_{1},v_{2},v_{3},v_{4},v_{5}}\}$ where $G[\{{v_{2},v_{3},v_{4},v_{5}}\}]\cong P_{4}$ and $v_{1}$ is a universal vertex. Consider $B\setminus\{{v_{i}}\}$ if $i=2,3,4,5$ then $B\setminus\{v_{i}\}$ contains $B_{1}$ and is thus a zero forcing set. Similarly, $B\setminus\{v_{1}\}$ contains $B_{2}$, so it is a zero forcing set. Therefore $B$ is a fault tolerant zero forcing set and $\operatorname{Z}_{t}(G)=5$. ∎
4 Compatible collections of path covers and applications to trees
4.1 Compatible path covers
The path cover number of a graph provides an important upper bound on the zero forcing number of the graph, and in this section we derive the corresponding parameter for fault tolerant zero forcing.
Definition 4.2.
A path cover of a graph $G$ is a collection of induced paths $\{P_{i}\}_{i=1}^{r}$ such that every vertex of $G$ lies in exactly one path $P_{i}$. The path cover number of a graph $G$, denoted $\mathrm{P}(G)$, is the minimum number of paths in a path cover of $G$.
Given a zero forcing set $B$ of a graph $G$, during the forcing process it may be possible that more than one blue vertex is able to force the same white vertex. Thus while the final set of blue vertices is unique for each initially blue set (whether it a zero forcing set or not), the forces used to obtain this final set of blue of vertices may not be. It is often convenient to work with a fixed set of forces, which leads to the next (well-known) definition.
Definition 4.3.
Let $G$ be a graph and $B$ a zero forcing set of $G$. A sequence of forces $v_{i}\rightarrow v_{i+1}$ listed in the order in which they occur is called a chronological list of forces $\mathcal{F}$ for $B$ in $G$. We omit $B,G$ when they are clear from the context, and denote by $\mathcal{F}\setminus u\rightarrow v$ the list of forces obtained by deleting $u\rightarrow v$ from $\mathcal{F}$.
Given zero forcing set $B$ of a graph $G$, a chronological list of forces $\mathcal{F}$ for $B$ in $G$ specifies a path cover of $G$. Each path in the path cover is of the form $G[\{v_{1},v_{2},\dots,v_{r}\}]$, where $v_{i}\rightarrow v_{i+1}\in\mathcal{F}$, $v_{1}\in B$ and $v_{r}$ does not perform a force in $\mathcal{F}$. These are induced paths since each of the forces in $\mathcal{F}$ is valid. In particular, this is true for minimum zero forcing sets, which leads to the following lower bound on the zero forcing number of a graph.
Theorem 4.4 ([22], Theorem 2.13).
Let $G$ be a graph. Then $\mathrm{P}(G)\leq\operatorname{Z}(G)$.
If $B$ is a fault tolerant zero forcing set, however, each $B\setminus\{v\}$ where $v\in B$ is a zero forcing set, and thus each vertex of $B$ induces a path cover of $G$. This suggests that collections of path covers are the natural object to consider in the fault tolerant case. However, a number of subtleties must be addressed before we define the relevant parameter.
Definition 4.5.
Let $G$ be a graph and $\mathcal{P}=\{P_{i}\}_{i=1}^{r}$ a path cover of $G$. Define $L(\mathcal{P})$ to be the endpoints of paths in $\mathcal{P}$, that is $L(\mathcal{P})=\{v\in V(T)\ |\ \deg_{\mathcal{P}}(v)\leq 1\}$. If $\mathfrak{P}$ is a set of path covers, then $L(\mathfrak{P})=\bigcup_{\mathcal{P}\in\mathfrak{P}}L(\mathcal{P})$.
Definition 4.6.
Let $G$ be a graph and $v\in V(G)$. A set $\mathfrak{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\dots,\mathcal{P}_{r}\}$ of path covers of $G$ is called compatible for $S\subseteq L(\mathfrak{P})$ if for every $v\in S$ both of the following hold:
  1. 1.
    $\deg_{\mathcal{P}_{i}}(v)\geq 1$ for all $i\in\{1,2,\dots,r\}$, and
  2. 2.
    for each $u_{i}\in L(\mathcal{P}_{i})$ such that $\deg_{\mathcal{P}_{i}}(v)=1$, there exists $\mathcal{P}_{j}$ such that $\deg_{\mathcal{P}_{j}}(v)=2$ or $\mathrm{comp}_{\mathcal{P}_{j}}(u_{i})\neq\mathrm{comp}_{\mathcal{P}_{j}}(v)$ and $\deg_{\mathcal{P}_{j}}(u_{i})\geq 1$.
While the definition of a compatible collection is quite technical, the idea is that if a collection is compatible for $S$, then the vertices of $S$ can always be forced from two different path covers and is not required to initiate forcing on path. Thus $S$ is not required in a minimal fault tolerant zero forcing set. We begin with a trivial example to show how compatible collections generalize path covers, before exhibiting a more interesting case.
Example 4.7.
Consider the graph $G$ in Figure 2.2. Since $G$ is a path, trivially $\mathcal{P}=\{G\}$ is path cover of $G$ and $\mathfrak{P}=\{\mathcal{P}\}$ is a compatible collection for each vertex in $\{v_{2},v_{3},\dots,v_{5}\}$ since $\deg(v_{i})=2$ for $2\leq i\leq 5$. However, $\deg(v_{1})=\deg(v_{7})=1$ and $v_{1}$ and $v_{7}$ are in the same component for every $\mathcal{P}\in\mathfrak{P}$, so $\mathfrak{P}$ is not compatible for $v_{1}$ and $v_{7}$.
8811223344556677
Figure 4.1: The graph $E_{1}$ in Example 4.8.
Example 4.8.
Consider the Graph $E_{1}$ in Figure 4.1. The collection $\mathfrak{P}=\{\mathcal{P}_{1},\mathcal{P}_{2},\mathcal{P}_{3},\mathcal{P}_{4}\}$, where each $\mathcal{P}_{i}$ is given in Figure 4.2, is a set of path covers compatible for $\{1,3,4,5\}$.
$\mathcal{P}_{1}8811223344556677$$\mathcal{P}_{2}$$\mathcal{P}_{3}8811223344556677$$\mathcal{P}_{4}$
Figure 4.2: A minimum compatible collection for $E_{1}$ in Example 4.8.
Definition 4.9.
Given a compatible collection of path covers $\mathfrak{P}$ for $S$, the irredundant vertices of $\mathfrak{P}$ are $I(\mathfrak{P})=L(\mathfrak{P})\setminus S$.
A collection of path covers $\mathfrak{P}$ of a graph $G$ is a minimum compatible collection for $G$ if $|I(\mathfrak{P})|\leq|I(\mathfrak{Q})|$ for every collection of path covers $\mathfrak{Q}$ of $G$. The cardinality of a minimum compatible collection for $G$ is called the irredundance number of $G$ is denoted by $\mathrm{I}(G)$.
Our next result shows that the irredundance number of a graph plays the same role in fault tolerant zero forcing that the path cover number plays in standard zero forcing.
Theorem 4.10.
If $G$ is a graph for which $\operatorname{Z}_{t}(G)$ exists, then $\mathrm{I}(G)\leq\operatorname{Z}_{t}(G)$.
Proof.
Let $B$ be a minimum fault tolerant set of $G$, then for each $v\in B$, $B\setminus\{v\}$ is a zero forcing set of $G$. Let $B^{\prime}_{v}$ be a minimal zero forcing set of $G$ such that $B^{\prime}_{v}\subseteq B\setminus\{v\}$. Then $B^{\prime}_{v}$ yields a path cover $\mathcal{P}_{v}$ of $G$. Let $\mathfrak{P}=\{\mathcal{P}_{v}\}_{v\in B}$. We claim $\mathfrak{P}$ is a compatible collection for each $w\not\in B$.
Consider $w\not\in B$. Since $w$ must be forced, we have $\deg_{\mathcal{P}_{v}}(w)\geq 1$ for every $v\in B$. If $w\not\in L(\mathfrak{P})$, then $\deg_{\mathcal{P}_{v}}(w)\geq 2$ for all $v\in B$ and $\mathfrak{P}$ is compatible for $w$. So suppose $w\in L(\mathfrak{P})$. For every $\mathcal{P}_{v}$ such that $\deg_{\mathcal{P}_{v}}(w)=1$, the other endpoint of $\mathrm{comp}_{\mathcal{P}_{v}}(w)$ must be in $B\setminus\{v\}$, and thus $\mathfrak{P}$ is compatible for $w$. Therefore $\mathfrak{P}$ is compatible for every $w\not\in B$, and since $I(\mathfrak{P})=|B|$, we have the result. ∎
Returning to Example 4.8, we see that $I(\mathfrak{P})=\{1,2,7,8\}$. This is indeed a fault tolerant zero forcing set, and it is minimal by Theorem 4.10 (or 2.5 and the fact that $\operatorname{Z}(E_{1})=3$).
In general, determining whether a set of compatible path covers for a graph $G$ is minimum is difficult and is a topic that requires further study. All path covers can be enumerated and thus an algorithm exists to test minimality, but this is not efficient to implement. Some simplifications exist, however, and as an example we offer the next result.
Lemma 4.11.
If $\mathfrak{P}$ is a minimum compatible collection for $G$ and $v\in L(\mathfrak{P})$ is isolated in every path cover of $\mathfrak{P}$, then $\mathfrak{P}-v$ is a minimum compatible collection of GvG-v. In particular, under these hypotheses I(G)=I(Gv)+1I(G)=I(G-v)+1
Proof.
Suppose to the contrary that $\mathfrak{Q}$ is a compatible collection for GvG-v and $I(\mathfrak{Q})<I(\mathfrak{P})-1$. Since $v$ was irredundant in $\mathfrak{P}$, adding $v$ to each path cover of $\mathfrak{Q}$ yields a compatible collection for $G$ with cardinality less than $I(\mathfrak{P})$, a contradiction. ∎
4.12 Trees
Trees are connected acyclic graphs, and we previously dealt with the special cases of $P_{n}$ and $K_{1,n}$. In this section we give an easily computable bound on the fault tolerant zero forcing number of a tree $T$, then show that the sets of compatible path covers of defined earlier in this section determine $\operatorname{Z}_{t}(T)$.
When we restrict our attention to trees, the inequality of Theorem 4.10 becomes equality, and establishing this is the primary goal of this section. We first recall the corresponding result in the standard zero forcing case.
Theorem 4.13 ([7]).
Let $T$ be a tree. Then $\mathrm{P}(T)=\operatorname{Z}(T)$.
In fact, there is a correspondence between (minimum) zero forcing sets of a tree $T$ and (minimum) path covers of $T$, obtained by taking one endpoint of each path in a path cover. This yields a zero forcing set of $T$, and all zero forcing sets are obtained in this way (see, for example, the proof of Proposition 4.2 in [2]).
We denote by $\ell(T)$ the number of vertices of degree 1 in a tree $T$, and refer to $\ell(T)$ as the leaf number of $T$. For leaky forcing, it was shown in [3, Proposition 3.2] that the set of leaves of a tree with an edge is the unique 1-leaky forcing set of $T$. This implies the next result, but we provide our own proof for self containment.
Proposition 4.14.
The set of leaves of a tree $T\not\cong K_{1}$ is a fault tolerant set of $T$. Hence,
$\operatorname{Z}_{t}(T)\leq\ell(t)$.
Proof.
We know from [18] that every tree has a zero forcing set that consists entirely of leaves. We must show that any set of all but one leaf is a zero forcing set of $T$.
For trees, a zero forcing set is obtained by choosing one endpoint of every path in a path cover of $T$. Let $B=\{v\in V(T)\ |\ v\text{ is a leaf of }T\}$. We consider $B\setminus\{v\}$ for some $v\in B$ and construct a path cover such that each path contains an endpoint in $B\setminus\{v\}=\{w_{1},w_{2},\dots,w_{r}\}$. Note that $r\geq 1$ since $T\not\cong K_{1}$. Furthermore, since $T$ is a tree there is unique path between any two vertices of $T$. Let ${P}_{1}$ be the unique path between $w_{1}$ and $v$. For $j\in\{2,3,\dots,r\}$, let $P_{j}$ be the maximum induced subpath of the unique path from $w_{j}$ to $v$ that $V(P_{j})\cap(\cup_{i=1}^{j-1}V(P_{i}))=\emptyset$. Clearly each path in $\mathcal{P}=\{P_{1},P_{2},\dots,P_{r}\}$ contains an endpoint. To see that $\mathcal{P}$ is a path cover of $T$, suppose $u\in V(T)$ s.t. $u\not\in V(P_{i})$ for every $i\in\{1,2,\dots,r\}$. Then $u$ cannot be a leaf, so $\deg_{T}(u)\geq 2$. Consider $x,y\in N_{T}(u)$ such that $x\neq y$, then there exists a unique path $Q_{1}$ to $v$ and a unique path $Q_{2}$ from $y$ to $v$. Then $T[V(Q_{1})\cup V(Q_{2})\cup\{u\}]$ is a cycle, a contradiction. Therefore $\mathcal{P}$ is a path cover of $T$ where every path contains at least one vertex of $B\setminus\{v\}$ as an endpoint. Thus $B\setminus\{v\}$ is a zero forcing set.
The star $K_{1,n}$ and the path $P_{n}$ show that the bound is tight. Inequality can be achieved, however, as the next example shows.
Example 4.15.
Let $G$ be the graph depicted in Figure 4.3. There are five leaves and thus $\ell(G)=5$, but $\operatorname{Z}_{t}(G)\leq 4$, as evidenced by the initially blue set $B=\{1,2,6,7\}$.
1122334455667788
Figure 4.3: The tree $T$ in Example 4.15 and Example 4.17.
Theorem 4.16.
Let $T$ be a tree with at least two vertices. Then $\operatorname{Z}_{t}(T)=|I(T)|$.
Proof.
In light of Theorem 4.10, we need only show that $I(T)\geq\operatorname{Z}_{t}(T)$. Thus, let $T$ be a tree and $\mathfrak{P}=\{\mathcal{P}_{i}\}_{i=1}^{r}$ a minimum compatible collection for $T$. Let $B=I(\mathfrak{P})$. We claim that $B$ is a fault tolerant zero forcing set of $T$, and proceed by induction on the order of $T$. When the order of $T$ is 2, we have $T=K_{2}$. There are two path covers of $K_{2}$ with the same endpoints, and thus we have $I(\mathfrak{P})=V(K_{2})$ which is clearly a fault tolerant set. Now assume the result holds for trees of order up to |T|1|T|-1.
For some $v\in B$, we consider $B\setminus\{v\}$. As stated previously, it is well-known that taking one endpoint of any path cover of $T$ yields a zero forcing set of $T$. Therefore if $\deg_{\mathcal{P}_{i}}(v)=1$ for some $i\in\{1,2,\dots,r\}$, then $B\setminus\{v\}$ contains an endpoint of every path in $\mathcal{P}_{i}$ and is thus a zero forcing set of $T$. Now suppose $v$ is isolated in every path cover $\mathcal{P}_{i}\in\mathfrak{P}$.
Let $w$ be a neighbor of $v$ in $T$. If $w\in B$ (and therefore $B\setminus\{v\}$) then for each $\mathcal{P}_{i}$, $w$ is either isolated or contained in a component with both endpoints. If $w$ is isolated in some $\mathcal{P}_{i}$, then since $B\setminus\{v\}$ is a zero forcing set of TvT-v, $w$ is able to force $v$ after the color change rule is applied to TvT-v, so $B\setminus\{v\}$ is a zero forcing set of $T$. Now suppose $w$ is never isolated in $\mathfrak{P}$. Then $\deg_{\mathcal{P}_{i}}(w)\geq 1$ for all $i$. If $\deg_{\mathcal{P}_{i}}(w)=2$ for all $i$ then $w\not\in I(\mathfrak{P})$ as hypothesized, thus there exists some $i$ such that $\deg_{\mathcal{P}_{i}}(w)=1$, and the other endpoint of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$ is in $B\setminus\{v\}$. Then as before, $w$ can force $v$ after all forces are carried out in TvT-v.
If $w\not\in B$, then $\mathfrak{P}$ is compatible for $w$. If there exists some $i$ such that $\deg_{\mathcal{P}_{i}}(w)=1$, and the other endpoint of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$ is in $B\setminus\{v\}$. So suppose $\deg_{\mathcal{P}_{i}}(w)=2$ for all $i$. Then both endpoints of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$ are in $B\setminus\{v\}$, so we can choose a forcing starting from one endpoint of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$ and terminating at $w$, while the other endpoint initiates a forcing process coloring the remainder of $\mathrm{comp}_{\mathcal{P}_{i}}(w)$. Then as before, $w$ can force $v$ after all forces are carried out in TvT-v.
Thus we have shown that $B$ is a fault tolerant zero forcing set, and therefore $\operatorname{Z}_{t}(T)\leq I(T)$.
Example 4.17.
Let $T$ be the tree in Figure 4.3. Then $\mathfrak{P}=\{\mathcal{P}_{1},\mathcal{P}_{2}\}$, where $\mathcal{P}_{1}=\{T[\{1,3,4,8\}],\\ T[\{5,6,7\}],T[\{2\}]\}$ and $\mathcal{P}_{2}=\{T[\{1,2,3\}],T[\{4,5,6,8\}],T[\{7\}]\}$ (shown in Figure 4.4) is a set of compatible path covers for the vertices $\{3,4,5,8\}$ in $T$. Furthermore, $\mathfrak{P}$ is a minimum compatible collection for $T$, and thus $I(\mathfrak{P})=\{1,2,6,7\}$ is a minimum fault tolerant set of $T$ and $\operatorname{Z}_{t}(T)=4$.
$1122334455667788$
Figure 4.4: Two path covers of the tree $T$ in Example 4.15.
5 Graph Operations
We now consider the effect of deleting a vertex or an edge from a graph on the fault tolerant zero forcing number. For a graph $G$ and $e$ an edge of $G$, denote by GeG-e the graph obtained by deleting $e$ from $G$ and by $G/\penalty 50e$ the graph obtained by contracting along $e$. For a vertex $v$ of $G$, denote by GvG-v the graph obtained by deleting $v$ and any incident edges to $v$ from $G$.
For zero forcing, this was studied in [14] and [24]. We begin by recalling the results therein.
Theorem 5.1 ([14],[24]).
Let $G$ be a graph, $e$ an edge of $G$, and $v$ a vertex of $G$.
  1. 1.
    $1\leq\operatorname{Z}(G)-\operatorname{Z}(G-e)\leq 1$
  2. 2.
    $1\leq\operatorname{Z}(G)-\operatorname{Z}(G/\penalty 50e)\leq 1$
  3. 3.
    $1\leq\operatorname{Z}(G)-\operatorname{Z}(G-v)\leq 1$
The proof of these results relies on Theorem 5.2, which is no longer true for fault tolerant zero forcing.
Theorem 5.2 ([6]).
Let $G$ be a graph. If $G$ is connected and has at least two vertices, then no vertex is in every minimal zero forcing set.
For example, we will show that all twins are required in every fault tolerant zero forcing set in Theorem 5.8. Thus, looser bounds hold in the fault tolerant case.
Definition 5.3.
Let $G$ be a graph and $B$ a zero forcing set of $G$. A sequence of forces $v_{i}\rightarrow v_{i+1}$ listed in the order in which they occur is called a chronological list of forces $\mathcal{F}$ for $B$ in $G$. We omit $B,G$ when they are clear from the context, and denote by $\mathcal{F}\setminus u\rightarrow v$ the list of forces obtained by deleting $u\rightarrow v$ from $\mathcal{F}$.
Theorem 5.4.
Let $G$ be a graph such that $\operatorname{Z}_{t}(G)$ exists and $e$ an edge of $G$ whose deletion leaves no isolated vertices. Then
$2\leq\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G-e)\leq 2.$
Proof.
Suppose $G$ and $e$ satisfy the hypotheses, let $u$ and $v$ be the endpoints of $e$, let $B$ be a fault tolerant zero forcing set of GeG-e and $\mathcal{F}_{w}$ be a chronological list of forces for $B\setminus\{w\}$ in GeG-e.
Let $B^{\prime}=B\cup\{u,v\}$. For each $\mathcal{F}_{w}$, construct $\mathcal{F}_{w}^{\prime}$ by removing any force received by $u$ or $v$. Each force in $\mathcal{F}_{w}^{\prime}$ is valid in $G$ since at each time step the number of white neighbors of each vertex of $G$ cannot be higher than the number of white neighbors at that time step in GeG-e. Thus $B^{\prime}$ is a fault tolerant zero forcing set for $G$, and $|B^{\prime}|\leq|B|+2$. Thus $\operatorname{Z}_{t}(G)\leq\operatorname{Z}_{t}(G-e)+2$.
Now suppose $B$ is a fault tolerant zero forcing set of $G$ with chronological list of forces $\mathcal{F}_{w}$ for $B\setminus\{w\}$. Construct $B^{\prime}=B\cup\{u,v\}$ and $\mathcal{F}_{w}^{\prime}$ from $\mathcal{F}_{w}$ by deleting any force received by $u$ or $v$. As before, each force in $\mathcal{F}_{w}^{\prime}$ is valid in GeG-e, so $B^{\prime}$ is a fault tolerant zero forcing set of GeG-e. Since $|B^{\prime}|\leq|B|+2$, we have $\operatorname{Z}_{t}(G-e)\leq\operatorname{Z}_{t}(G)+2$.
112233445566778899aabbeeddcc
Figure 5.1: The graph in Example 5.5.
Example 5.5.
To show that all of the possibilities of the value of $\operatorname{Z}_{t}(G)$ and $\operatorname{Z}_{t}(G-e)$ implied by Theorem 5.4 can be realized, consider the graph $G$ in Figure 5.1. Note that no graph of order less than 9 exhibits all possibilities. Computation of $\operatorname{Z}_{t}(G)=4$ and $\operatorname{Z}_{t}(G-e)$ for various edges $e$ establishes the following:
To show that all of the possibilities of the value of $\operatorname{Z}_{t}(G)$ and $\operatorname{Z}_{t}(G-e)$ implied by Theorem 5.4 can be realized, consider the graph $G$ in Figure 5.1. Note that no graph of order less than 9 exhibits all possibilities.8 Computation of $\operatorname{Z}_{t}(G)=4$ and $\operatorname{Z}_{t}(G-e)$ for various edges $e$ establishes the following:
$\mathbf{e}$ $\mathbf{\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G-e)}$
a 2
b 1
c 0
d -1
e -2
Theorem 5.6.
Let $G$ be a graph such that $\operatorname{Z}_{t}(G)$ exists, and $e$ an edge of $G$ such that $\operatorname{Z}_{t}(G/\penalty 50e)$ exists. Then
$2\leq\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G/\penalty 50e)\leq 2.$
Proof.
Suppose $u$ and $v$ are the endpoints of $e$, denote the quotient map from $G$ to $G/\penalty 50e$ by $\varphi$, and let $u^{\prime}=\varphi(u)=\varphi(v)$. Let $B$ be a minimum fault-tolerant zero-forcing set of $G/\penalty 50e$. Suppose first $u^{\prime}\not\in B$. Let $F$ be a chronological list of forces for $B\setminus\{{w}\}$ in $G/\penalty 50e$ for some $w\in B$. Then each force in $F$ is valid until $x\rightarrow u^{\prime}$ for some $x\in V(G/\penalty 50e)$. Let $F^{\prime}$ be the chronological list of forces that is identical to $F$ until this force. After applying the forces in $F^{\prime}$, $x$ has at most has two white neighbors in $G$, namely $u$ and $v$.
Note that since $u^{\prime}\not\in B$, $\varphi$ is bijective on $B$. Let $D=\varphi^{-1}(B)\cup\{{u,v}\}$. Consider $D\setminus\{{w}\}$ for some $w\in D$. After performing all the forces in $F^{\prime}$, $x$ has at most one white neighbor. If $x$ has exactly one white neighbor, without loss of generality $u$, add $x\rightarrow u$ to $F^{\prime}$.
If $u^{\prime}\rightarrow y\in\mathcal{F}$, for some $y\in V(G/e)$, then $y$ is the only white vertex in $N(u)\cup N(v)$. So $u\rightarrow y$ or $v\rightarrow y$ is valid. Add whichever force is valid to $\mathcal{F}^{\prime}$, and the rest of the forces after $u^{\prime}\rightarrow y$ in $\mathcal{F}$ are valid in $G$, so we append them to $\mathcal{F}^{\prime}$. Thus $D\setminus\{{w}\}$ is a zero forcing set for $G$ with chronological list of force $\mathcal{F}^{\prime}$. Since $w$ was arbitrary, $D$ is a fault tolerant zero forcing set of $G$.
Now consider the case where $u^{\prime}\in B$. Let $\mathcal{F}$ be a chronological list of forces for $B\setminus\{w\}$ in $G/\penalty 50e$ where $w\neq u^{\prime}$. If $u^{\prime}\rightarrow c\in\mathcal{F}$ and occurs at position $j$ in $\mathcal{F}$, then define $B^{\prime}=\varphi^{-1}(B\setminus\{u^{\prime}\})\cup\{u,v,c\}$, or $B^{\prime}=\varphi^{-1}(B\setminus\{u^{\prime}\})\cup\{u,v\}$ if $u^{\prime}$ does not perform a force in $\mathcal{F}$.
Consider $B^{\prime}\setminus\{r\}$. If $\{r\}\cap\{u,v\}=\emptyset$, then since $N_{G/\penalty 50e}(u^{\prime})=N_{G}(u)\cup N_{G}(v)$ replace $u^{\prime}\rightarrow c$ if it exists with $u\rightarrow c$ or $v\rightarrow c$ at position $j$ (at least one of which must be valid).
Consider $B^{\prime}\setminus\{r\}$. If $\{r\}\cap\{u,v\}\neq\emptyset$, then without loss of generality suppose $\{r\}\cap\{u,v\}=\{u\}$. $N(u)\cup N(v)$ has exactly one white vertex, namely $u$. Replace $u^{\prime}\rightarrow c$, if it exists, with $v\rightarrow u$ at position $j$. Since $|B^{\prime}|\leq|B|+2$, then $\operatorname{Z}_{t}(G)\leq\operatorname{Z}_{t}(G/\penalty 50e)+2$.
For the other inequality, let $B$ be a minimum fault tolerant zero forcing set for $G$. We will construct a fault tolerant zero forcing set $B^{\prime}\supset\varphi(B)$. Let $\mathcal{F}$ be a chronological list of forces for $B\setminus\{{w}\}$ in $G$, where $w\in B$. Each force $a\rightarrow b\in\mathcal{F}$ is valid in $G/e$ as long as $\{{a,b}\}\cap\{{u,v}\}=\emptyset$. Let $\mathcal{F}^{\prime}$ be the chronological list of forces that is identical to $\mathcal{F}$ except for the deletion of forces of this form.
There are several cases to consider. Suppose first that $a\rightarrow u\in\mathcal{F}$, where $a\neq v$. Since $a$ has one white neighbor in $G$, $a$ has at most one white neighbor in $G/e$, namely $u^{\prime}$. In this case, add $a\rightarrow u^{\prime}$ to $\mathcal{F}^{\prime}$ in the location where $a\rightarrow u$ was in $\mathcal{F}$.
Now suppose $u\rightarrow v\in\mathcal{F}$, then after the preceding forces in $\mathcal{F}$ occur, $u$ is blue. Then at this point in $\mathcal{F}^{\prime}$, $u^{\prime}$ is also blue. Suppose $u\rightarrow c,v\rightarrow d\in\mathcal{F}$. Let $B^{\prime}=\varphi(B)\cup\{{c,d}\}$ so that the remaining forces for $B\setminus\{{w}\}$ are all valid for $B^{\prime}$ in $G/\penalty 50e$ where $w\not\in\{c,d\}$. Consider $B^{\prime}\setminus\{{c}\}$ (the case of $B^{\prime}\setminus\{d\}$ is similar).
Every force occurring before $u\rightarrow c$ in $\mathcal{F}$ is valid in $G/\penalty 50e$. Furthermore, since $d$ is filled, $u^{\prime}$ has at most one white neighbor after these forces have been performed. Add $u^{\prime}\rightarrow c$ to $F^{\prime}$ in the location where $u\rightarrow c$ occurred in $\mathcal{F}$. Then $\mathcal{F}^{\prime}$ is a chronological list of forces of $B^{\prime}\setminus\{c\}$ in $G/\penalty 50e$, so $B^{\prime}\setminus\{c\}$ is a zero forcing of $G/\penalty 50e$. This resolves the remaining cases to show that $B^{\prime}$ is a fault tolerant zero forcing set of $G/\penalty 50e$, and thus $\operatorname{Z}_{t}(G/\penalty 50e)\leq\operatorname{Z}_{t}(G)+2$. ∎
1122334455667788(a)aaccbbdd11223344(b)$f$
Figure 5.2: The graphs in Example 5.7.
Example 5.7.
Consider the graph $G$ in Figure 5.2 (a). Computation of $\operatorname{Z}_{t}(G)$ and $\operatorname{Z}_{t}(G/\penalty 50e)$ for various edges $e$ establishes the following:
$e$ $\mathbf{\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G/\penalty 50e)}$
$a$ -2
$b$ -1
$c$ 0
$d$ 1
Meanwhile, for the graph $H$ in Figure 5.2 (b), we have $\operatorname{Z}_{t}(H)-\operatorname{Z}_{t}(H/\penalty 50f)=2$.
Theorem 5.8.
Let $G$ be a graph such that $\operatorname{Z}_{t}(G)$ exists and $v$ a vertex of $G$ whose deletion leaves no isolated vertices. Then
$\deg(v)\leq\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G-v)\leq 2.$
Proof.
Let $G$ and $v$ satisfy the hypotheses and let $B$ be a fault tolerant zero forcing set of GvG-v with corresponding chronological list of forces $\mathcal{F}_{w}$ for each $B\setminus\{w\}$. Let $v_{1}\in B$, then there is a first force $x\rightarrow w$ in $\mathcal{F}_{v_{1}}$. Construct $B^{\prime}=B\cup\{v,w\}$ and for each $u\in B$ we let $\mathcal{F}_{u}^{\prime}=\mathcal{F}_{u}\setminus y\rightarrow w$, where $y$ is the vertex of GvG-v that forces $w$. Each of these forces is valid and results in all vertices blue when starting with $B^{\prime}\setminus\{u\}$ initially blue.
It remains to show that $B^{\prime}\setminus\{v\}$ is a zero forcing set. Before any forces are performed, $v$ is the unique white neighbor of $x$, and therefore $x\rightarrow v$ is a valid first force. We then proceed with the forces in $\mathcal{F}^{\prime}_{v_{1}}$ to color the remaining vertices. Thus $B^{\prime}$ is a fault tolerant zero forcing set with $|B^{\prime}|\leq|B|+2$, so $\operatorname{Z}_{t}(G)\leq\operatorname{Z}_{t}(G-v)+2$.
For the other inequality, let $B$ be a minimum fault tolerant zero forcing set of $G$. Since $v$ can only force its neighbors, $B\cup N_{G}(v)$ is a fault tolerant zero forcing set of GvG-v. ∎
11223344 55
Figure 5.3: The graph in Example 5.10.
Theorem 5.9.
Let $G$ be a graph with twins $u$ and $v$. Then $0\leq\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G-v)\leq 2$.
Proof.
We have $\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G-v)\leq 2$ by Theorem 5.8, thus it remains to show that $\operatorname{Z}_{t}(G-v)\leq\operatorname{Z}_{t}(G)$. Let $B$ be a minimum fault tolerant zero forcing set of $G$, and $B$ contains both $u$ and $v$. Consider $B\setminus\{x\}$ for some $x\in B$ and a chronological list of forces $\mathcal{F}$ for $B\setminus\{v,x\}$. We create $\mathcal{F}^{\prime}$ from $\mathcal{F}$ by replacing a force of the form $v\rightarrow w$ with $u\rightarrow w$ if it appears, and if $x=v$, removing the force where $v$ is forced. Then $\mathcal{F}^{\prime}$ is a chronological list of forces for $B\setminus\{x,v\}$ in GvG-v that forces all vertices. Thus $B\setminus\{v\}$ is a fault tolerant zero forcing set. ∎
Note that by applying Theorem 2.11, $\operatorname{Z}_{t}(G-v)$ exists only if deleting $v$ does not result in any isolated vertices.
Example 5.10.
Consider the graph $G$ in Figure 5.3. Computation of $\operatorname{Z}_{t}(G)$ and $\operatorname{Z}_{t}(G-v)$ for various vertices $v$ establishes the following:
$\mathbf{v}$ $\mathbf{\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G-v)}$
5 0
1 1
2 2
There is not a useful lower bound on $\operatorname{Z}_{t}(G)-\operatorname{Z}_{t}(G-v)$, as the next example shows.
Example 5.11.
Let $G$ be a graph with $V(G)=\{v_{1},v_{2},\dots,v_{n}\}$ and $E(G)=\{v_{1}v_{i}\ |\ i=2,3,\dots,2n+1\}\cup\{v_{i}v_{i+1}\ |\ i\equiv 0\mod 2\}$. This graph is commonly called a generalized star. We can calculate $\operatorname{Z}_{t}(G)=n$, while $\operatorname{Z}_{t}(G-v_{1})=2n$.
We note that Theorem 5.9 could be used to shorten the proofs of Proposition 3.2 - Proposition 3.4. In the case of graphs with several twins, it also reduces the search space in brute-force calculations of the fault tolerant zero forcing number.